Suppose your firm operates a chain of video rental stores in South Bend. Your firm’s beta is 1.35, its cost of debt is 6.25% at its current D/E ratio of 0.30, and its marginal tax rate is 21%. Assume a risk-free rate of 3.2%, a market risk premium of 5.8%.

a. You are planning to expand through the purchase of a privately-held rival chain of video rental stores in the SB area. What discount rate should you use for your analysis of this expansion?

b. You are also considering acquiring a privately-held takeout pizza business. A comparable public firm (Pizza World) has a stock beta of 0.95, a D/E ratio of 0.20, and tax rate of 20%. If you will finance the acquisition with your usual mix of debt and equity, what cost of capital should you use for your analysis of this acquisition?
Answer #1

1.
=0.3/1.3*6.25%*(1-21%)+1/1.3*(3.2%+5.8%*1.35)=9.62403846153846%

2.
=0.3/1.3*6.25%*(1-21%)+1/1.3*(3.2%+5.8%*(0.95/(1+(1-21%)*0.2)*(1+(1-21%)*0.3)))=8.12857546167132%